OpenStudy (anonymous):

Simplify the expression: 14/(3x) + (x+5)/(3x)

OpenStudy (anonymous):

\[ \frac{14}{3x} + \frac{(x+5)}{3x}\] So what can you tell about the two terms' denominators?

OpenStudy (anonymous):

And also remember that, similarly, \[\frac{1}{2} + \frac{1}{2} = \frac{(1+1)}{2} = \frac{2}{2}=1\]

OpenStudy (anonymous):

so 1 is the answer? really

OpenStudy (anonymous):

No no no, sorry. That as just an example of two two with the same denominator. It has nothing to do calculation wise with the problem at hand.

OpenStudy (anonymous):

~Two terms

OpenStudy (anonymous):

so what is the answer?

OpenStudy (anonymous):

Have a go at it first. Just like in the example, they both have the same denominator (bottom term) - in your problem, it's 3x; so you add the numerators and put the sum over their common denominator.

OpenStudy (anonymous):

i really cant this isn't my homework its my other friends i told her i would see if could get the answer idk how do this haven't learned it yet shes older

OpenStudy (anonymous):

Well, \[14 + x + 5=?\]

OpenStudy (anonymous):

You can think of it as \[ x + (14+5) = x + \ ?\]

OpenStudy (anonymous):

What would the question mark be?

OpenStudy (anonymous):

(promise it'll take two seconds after this, and you can show your friend and impress them! Much better than just an answer.)

OpenStudy (anonymous):

ok thank you :)

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

Following through, the question mark would be 14+5 \[14+x+5 = x + (14+5) = x+19\] This is the sum of the numerator \[\frac{\overset{\text{this top term}}{14}}{3x}+\frac{\overset{\text{and this top term}}{(x+5)}}{3x}\] So then the answer would be that sum above (x+19) all divided by the denominator (3x) \[\frac{14}{3x} + \frac{(x+5)}{3x} = \frac{14+x+5}{3x}\] Follow?

OpenStudy (anonymous):

not really im only in 7th i have no idea what your doing she is in the 12 grade lol

OpenStudy (anonymous):

No worries ^_^ It's exactly like adding fractions, except with a few letters thrown in. ^_^ in any case, the answer your friend is looking or is \[\frac{x+19}{3x}\]

OpenStudy (anonymous):

thank you mamn

OpenStudy (anonymous):

Very welcome - happy start of the semester!